OpenStudy (anonymous):

Inspector A visually inspected 1000 ceramic mugs and found flaws in 36 of them. Inspector B inspected teh same 1000 ceramic mugs and found flaws in 42 of them. A total of 950 mugs were found to be good by both inspectors. one of the 1000 mugs is selected randomly. (a) find the probability that a flaw was found in this mug by at least one of the two inspectors. (b) find the probability that the flaws were found in this mug by both inspectors. (c) find the probability that a flaw was found by inspector b but not by inspector A.

OpenStudy (experimentx):

venn diagram ... start with venn diagram.

OpenStudy (anonymous):

so for (a) it is P(A U B) im not so sure about how to go with the venn diagram

OpenStudy (experimentx):

yep ... \[ P(A \cup B) + P((A \cup B)^c) = 1000 \\ P(A \cup B) = P(A) + P(B) - P(A \cap B)\]

OpenStudy (experimentx):

\[ P((A \cup B)^c) = 950 \]

OpenStudy (experimentx):

i think you can do the rest of it.

OpenStudy (anonymous):

im still lost.... so if P(AUB) ') = 950 then P(A U B)= 50?

OpenStudy (experimentx):

of course yes.

OpenStudy (anonymous):

so solving for P (A and B) = 49.92

OpenStudy (experimentx):

hell no!! use equation 2

OpenStudy (anonymous):

k thanks